14. In the following electrical circuit, if the voltage supply Vs 220 V (Calculate voltage across R5, current passing through R5, and Pes the power dissipated by R5. R3 R4 R5 Vs 220VT Ri 12Ω 100Ω
Answer #1

In the given circuit All resistor is connected directely parallel to the voltage supply, So voltage across R5 is equal to supply voltage.

i.e V_R5 = 220 V.

Current passing through R5

I5 = V_R5/ R5

I5 = 220/200 = 1.1 A.

Power dissipiated by R5

P_R5 = (I5)^2*R5 = (1.1)^2*200 = 242 W.

Hope this helps! Please thumbs up?.